Equation of the motion of a Body moving along x axis at an instant t second is given by x=40+12t+t^3 displacement of the particle before comming to rest is 

Grade:11

1 Answers

Arun
6 years ago
You have typed the wrong equation for x
Because in you equation velocity can never be zero.
Hence I have taken x= 40 + 12t - t^3
 
find the velocity and time relation by differentiating the given equation w.r.t t.
v=dx/dt =12 - 3 t^2. The body comes to rest when velocity v= 0.
12 -3 t^2 =0, t = 2 secs.
At t= 2 , x =40 + 12 *2 - 2 ^3= 56 m